Pete bought a new coffee machine for $6,000 to replace the old machine. Suppose the machine is estimated to have a life of 3 years and no residual value at the end of the life. What are the depreciation expenses over the 3-year period using the straight-line and MACRS methods (Please show work)

The 3-year depreciation schedule of the MACRS method is as follows.

Year 1          2    3 4   

%    33.33 44.45 14.81    7.41       

Year Straight Line Depreciation Expense MACRS Depreciation Expense
1
2
3
4
Total Depreciation

  

Art Angel operated a small seasonal lake marina, renting boats. He runs the business for only four months per year. He bought three new boats for $10,000 each. The new boats are estimated to have a 10-year life and a residual (trade-in) value of $1,200 each. Please answer the following questions. What are the depreciation expenses for the three boats per month using the straight-line method?
Answer #1

1.

Year Straight Line Depreciation Expense MACRS Depreciation Expense
1 $ 2,000 $ 1,999.80
2 2,000 2,667
3 2,000 888.60
4 0
Total Depreciation $ 6,000 $ 5,555.40

Straight-line depreciation expense per year = ( Cost - Residual Value ) / Estimated Useful Life = $ ( 6,000 - 0 ) / 3 = $ 2,000.

MACRS depreciation expense, Year 1 = $ 6,000 x 33.33 % = $ 1,999.80

MACRS depreciation expense, Year 2 = $ 6,000 x 44.45 % = $ 2,667

MACRS depreciation expense, Year 3 = $ 6,000 x 14.81 % = $ 888.60.

2. Total cost of the boats = $ 10,000 x 3 = $ 30,000.

Estimated residual value = $ 1,200 x 3 = $ 3,600.

Straight Line depreciation expense per year = ( Cost - Residual Value ) / Estimated Useful Life = $ ( 30,000 - 3,600) / 10 = $ 2,640.

Straight Line depreciation expense per month = $ 2,640 / 12 = $ 220